Find the simple interest and the total payback amount on a loan of $600 at an annual rate of 8% for 4 months.

Answer by checkley75(3666)

600(.08*4/12)=600(.32/12)=192/12=$16 IS THE SIMPLE INTEREST. THE TOTAL PAYBACK=600+16=$616.